How do i solve the system 3x+2y=4 and y-2x=9
Found 2 solutions by richard1234, josh_jordan:
Answer by richard1234(7193)

y-2x = 9 <--> y = 2x+9 from second eq.

Substitute this expression for y into the first equation and solve for x. Then solve for y.

Answer by josh_jordan(263)

The easiest way to solve this system of linear equations is by substitution, since one of the variables has no number in front of it. Since the second equation contains a letter without a number in front of it, we will isolate that letter on one side of the equal sign, so y - 2x = 9 will become:

y = 9 + 2x

Next, we will replace y in the first equation, with 9 + 2x:

3x + 2(9 + 2x) = 4

Now, multiply 2 by everything inside the parenthesis, which will give us the following equation:

3x + 18 + 4x = 4

Combine 3x and 4x, which gives us:

7x + 18 = 4

Subtract 18 from both sides of the equal sign, giving us:

7x = -14

Divide both sides of the equal sign by 7, in order to get x by itself. This gives us:

x = -2

Now that we have the value of x, we can replace x with -2 in our rewritten second equation: y = 9 + 2x will become:

y = 9 + 2(-2), which gives us:

y = 9 - 4, which leaves us with

y = 5

Therefore, x = -2 and y = 5